# Non-metricity approach to Jackiw-Teitelboim gravity

### v1

@@ I Introduction
In general, we may construct a gravity theory based on the gauge theory of the Poincaré algebra. The Poincaré algebra is composed of the generators of the Lorentz transformation and the translation. By requiring the curvature of the translation to vanish, the curvature of the Lorentz transformation gives the curvature of Einstein gravity. The curvature of the translation corresponds to the torsion tensor. In [16], the action is chosen to be a square of the curvature corresponding to the translation, that is, field strength, which is an analogue of the electrodynamics or non-abelian gauge theory.
@@
In this paper, we construct JT gravity in the framework of symmetric teleparallel gravity using a non-metricity tensor [21, 22, 23, 24], where both the curvature and the torsion vanish. In symmetric teleparallel gravity, we often use the scalar quantity $Q$ composed of quantities called non-metricity tensors. From the viewpoint of covariance, $Q$ is not a unique scalar quantity although the difference between $Q$ and the scalar curvature in Einstein gravity is a total derivative as in the case of the torsion scalar. If we consider the perturbation in the flat Minkowski background, $Q$ reduces to the Lagrangian density of the gravitational wave. This shows that $Q$ could be a unique scalar quantity composed of bilinear of the non-metricity tensors so that the action linear to $Q$ gives only the massless spin-two mode but does not include extra modes, which are often ghosts. In two dimensions, the situation is rather special because there is no propagating mode in Einstein gravity. Therefore if one considers more general scalar quantities composed of the non-metricity tensors, there could be any propagating modes. Such modes are often ghosts, whose kinetic energy is not bounded below. If we change the signature of the action, the mode becomes canonical and the kinetic energy is bounded below. In the higher dimensions, if we change the signature of the action, the standard gravitational wave becomes a ghost and the kinetic energy becomes unbounded below. In the two dimensions, because the propagating massless spin-two mode does not exist, even if we change the signature of the action, there might not appear any ghost.
@@
A problem proper to the symmetric teleparallel gravity is the treatment of the connections. In symmetric teleparallel gravity, the connections are degrees of freedom independent from metric although the connections are constrained by conditions that require the curvature and the torsion to vanish. The structure of the constraints is very complicated and the number of degrees of freedom is not clear even off-shell. Therefore it is very difficult to count the number of physical degrees of freedom. We often use the coincident gauge condition where the connections vanish. This gauge condition makes the equations simple but the structure of the constraints is still not clear and because the coincident gauge fixes the general covariance, the solutions severely depend on the coordinate choice and it becomes difficult to find the coordinate system consistent with the coincident gauge. In order to clarify the off-shell degrees of freedom, in [25], the metric and four scalar fields are chosen to be independent fields off-shell. Although the number of physical degrees of freedom is still not clear, all the equations are given in a closed form.

### v2

@@ I Introduction
In general, we may construct a gravity theory based on the gauge theory of the Poincaré algebra. The Poincaré algebra is composed of the generators of the Lorentz transformation and the translation. By requiring the curvature of the translation to vanish, the curvature of the Lorentz transformation gives the curvature of Einstein’s gravity. The curvature of the translation corresponds to the torsion tensor. In [16], the action is chosen to be a square of the curvature corresponding to the translation, that is, field strength, which is an analogue of the electrodynamics or non-abelian gauge theory.
@@
In this paper, we construct JT gravity in the framework of symmetric teleparallel gravity using a non-metricity tensor [21, 22, 23, 24], where both the curvature and the torsion vanish. In symmetric teleparallel gravity, we often use the scalar quantity $Q$ composed of quantities called non-metricity tensors. From the viewpoint of covariance, $Q$ is not a unique scalar quantity although the difference between $Q$ and the scalar curvature in Einstein’s gravity is a total derivative as in the case of the torsion scalar. If we consider the perturbation in the flat Minkowski background, $Q$ reduces to the Lagrangian density of the gravitational wave. This shows that $Q$ could be a unique scalar quantity composed of bilinear of the non-metricity tensors so that the action linear to $Q$ gives only the massless spin-two mode but does not include extra modes, which are often ghosts. In two dimensions, the situation is rather special because there is no propagating mode in Einstein’s gravity. Therefore if one considers more general scalar quantities composed of the non-metricity tensors, there could be any propagating modes. Such modes are often ghosts, whose kinetic energy is not bounded below. If we change the signature of the action, the mode becomes canonical and the kinetic energy is bounded below. In the higher dimensions, if we change the signature of the action, the standard gravitational wave becomes a ghost and the kinetic energy becomes unbounded below. In the two dimensions, because the propagating massless spin-two mode does not exist, even if we change the signature of the action, there might not appear any ghost. We should note that the symmetric teleparallel gravity in two dimensions was first considered in [25] and the general scalar combinations composed of the non-metricity tensors were also investigated in [26, 27].
@@
A problem proper to the symmetric teleparallel gravity is the treatment of the connections. In symmetric teleparallel gravity, the connections are degrees of freedom independent from metric although the connections are constrained by conditions that require the curvature and the torsion to vanish. The structure of the constraints is very complicated and the number of degrees of freedom is not clear even off-shell. Therefore it is very difficult to count the number of physical degrees of freedom. We often use the coincident gauge condition invented in [28, 29], where the connections vanish. This gauge condition makes the equations simple but the structure of the constraints is still not clear and because the coincident gauge fixes the general covariance, the solutions severely depend on the coordinate choice and it becomes difficult to find the coordinate system consistent with the coincident gauge. In order to clarify the off-shell degrees of freedom, in [30], the metric and four scalar fields are chosen to be independent fields off-shell. Although the number of physical degrees of freedom is still not clear, all the equations are given in a closed form.